# A general framework for the generation of probabilistic socioeconomic scenarios: Quantification of national food security risk with application to the cases of Egypt and Ethiopia

### 07 Dec 2023 v1

@@ 2.2 Shared Socioeconomic Pathways, Population and GDP Projections
One advantage of this methodology is that the corresponding levels for the scenarios are not preassigned, as for example in the UN methodology, but are determined endogenously by the history and dynamics of the data from the population model 999which by its Bayesian nature is calibrated using a vast global database of past population data, hence carries important information on the history of the population phenomenon.. Repeating the procedure above for each of the trajectories in the sample for each country, we end up with three sub-samples (one triplet for fertility rate and one triplet for life expectancy resp.) each corresponding to different possible realizations of the low, medium and high intensity levels (sub-scenarios). These sub-samples provide important statistical information, such as moments, variability, etc. within scenarios (i.e. within the Low, Medium and High sub-scenarios described in Table 1), and can be used to generate further scenarios and the calculation of conditional expected values of other quantities of interest, depending on the fundamental quantities (i.e. TFR or $e_{0}$ etc) modeled by these trajectories. A possible question to this procedure could be concerning the validity of the generated sample of trajectories by the Bayesian model for the key population drivers (TFR, $e_{0}$) and its ability to represent all possible (realistic) future states of the world. Since the Bayesian model relies on the observed data from previous periods and takes into account possible relations of the country or region under study with all other countries and regions of the world, then any reasonable scenario (with respect to the data that have been collected up to the time the projection task is executed) should be amenable to simulation. Using a sufficiently large number of simulated trajectories, e.g. 100,000 trajectories, should guarantee that the results are reliable. The country-specific discrimination rule for each population driver is illustrated in Table 2. Note that migration levels could be also determined in the same manner if a probabilistic approach had been used (see e.g. Azose and Raftery (2015); Azose et al. (2016)), however in our case for simplicity, and since for the application we consider migration will not play a major role we use the pointwise net migration projections under each SSP scenario, as provided by the Wittgenstein Center database.
@@
We follow this approach for each country of interest and generate a sufficient database of possible future realizations of the population phenomenon, as obtained by model (1), which would be compatible with the various SSP scenarios, and the associated sub-scenarios on the population key quantities. For example, for a country belonging to the HiFert group (according to Table 1), the SSP2 scenario consists of the TFR trajectories that belong to the Medium TFR sub-scenario, the e0 trajectories that belong to the Medium e0 sub-scenario and the net migration projection under the medium sub-scenario. These three components are integrated by the population model (1) and provide the population trajectories out of which the SSP2 scenario for this country consists of. This is done in a probabilistic manner, in the sense that we are able to compute quantiles, moments, etc. for the certain population dimensions for this country. In Figure 2 the resulting median population projections for Egypt (EGY) and Ethiopia (ETH) are illustrated under each SSP scenario. Moreover, the population projections in terms of relative growth rates including also the 90% uncertainty zone for both countries in each scenario are illustrated in Figure 10 in Appendix A, while the instantaneous projected total population distributions at year 2050 for Egypt and Ethiopia under all SSP scenarios are illustrated in Appendix B.
Building scenarios representing different pathways for the population and its age structure evolution, offers a vehicle for the estimation of key economic drivers, like labour force, gross domestic product (GDP) and others (see e.g. Aksoy et al. (2019)). In this paper such economic drivers scenarios (i.e. employing the various SSPs) are obtained using the global dynamic economic model MaGE Fouré et al. (2013); Fontagné et al. (2022). This model was developed by J. Fouré, A. Bénassy-Quéré and L. Fontagné in CEPII and is freely available from CEPII’s website [^10] http://www.cepii.fr/cepii/en/bdd_modele/presentation.asp?id=13 in Stata. MaGE assumes that the world consists of economies of individual countries with each country $c$ characterized at time $t$ by a three-factor CES (Constant Elasticity of Substitution) production function with the capital and labour contributions modelled by the Cobb-Douglas parametric form
@@
where $I$ denotes the GDP, $K$ the capital, $L$ the labour force111111child labour is not modelled by MaGE, $E$ the energy consumption with $c$ denoting the country and $t$ corresponding either to 1-year periods or 5-years periods. The elasticity parameters are assumed to be the same for all countries ($\alpha=0.31$, $\sigma=0.136$) and constant in time, while the parameters $A$ (Total Factor Productivity or TFP) and $B$ (Energy Productivity or Energy Efficiency121212where energy efficiency is defined as an analogous quantity to the ratio between energy consumption and GDP) are assumed to be country specific and temporally varying. The model depends on population primarily through labour force and secondarily through the life cycle savings modelling which is introduced in the modelling of investment. MaGE allows the user to run generalized scenarios concerning the future states of the world economy, compatible with the SSP scenarios as far as population is concerned and produce relevant projections for each SSP scenario. In particular the model included in MaGE allows for the integration of demographics (including education) and economics to allow for predictions concerning population growth and economic quantities such as economic growth rates, GDP, energy consumption, etc.
@@ 2.3 A Combined Framework with Climate Scenarios
Since we are investigating food security issues, critical environmental drivers such as temperature, precipitation, water stress, etc. must also be taken into account, given that they affect directly agricultural operations and thus the food system capacity. The so called Representative Concentration Pathways (RCPs) Van Vuuren et al. (2011) determine scenarios regarding the increase of the mean temperature of the planet by the time horizon 2100, taking into account whether certain environmental policies are adopted or not. These scenarios are characterised by the levels of the increase of the mean temperature of the planet and the intensity of the mitigation measures that have to be adopted per scenario (please see Table 7 in Appendix A for the specifications per RCP scenario).
@@
The modern trend in creating realistic scenarios for the future states of the world, is to combine the concept of SSP with that of RCP scenarios. At a first glance, one would naively provide 30 different SSP-RCP scenarios (given the ones illustrated in Table 1 and the standard six RCP scenarios). However, this is not exactly the case since the SSP and RCP scenarios, although referring to different target quantities, are not independent. The Coupled Model Intercomparison Project [^13] https://www.wcrp-climate.org/wgcm-cmip/wgcm-cmip6 (CMIP) has published interesting studies on these climate scenarios by integrating different environmental models and using very dense databases. The combined scenarios (SSP-RCP) that are up to now well tested and available are illustrated in Table 3, consisting of six different scenarios with varying levels of optimism concerning the phenomenon of temperature change Meinshausen et al. (2020). Thinking along the direction of providing food security risk141414in the sense discussed in Section 1 evaluations in a framework which is compatible with both the socioeconomic and climate pathways, we perform estimations according to the six scenarios illustrated in Table 3 building the relevant models for food capacity accordingly (see Section 3).
3 Probabilistic estimations and projections on national food security risk
Predicting future food shortages is of paramount importance to policy makers, as it allows for the adoption of proactive measures, such as for example the restructuring of food production, land use, etc or planning for imports. The design and adoption of effective policies require a deeper understanding of the underlying forces that drive potential major changes in the food system. In this section we propose a (minimal) statistical model, requiring as little as possible and well documented publicly available data, that is able to provide at country level, probabilistic projections of minimum caloric requirements and food system capacity, and hence quantifying food security risk. Population growth, which is the key driver in this model, is treated in a fully probabilistic fashion using the detailed model of Raftery et al. (2013) and its modifications introduced in Section 2. For the modelling task of food security risk, socioeconomic drivers (e.g. GDP, labour, etc), natural resources (e.g. water resources, cropland area, etc) and climate variables (e.g. temperature, precipitation) are involved, however, are not treated in a fully probabilistic fashion. This is partly due to the lack of sufficient computational resources, data or suitable models (such an endeavour would turn the resulting model extremely computationally demanding and also introduce difficulties in its interpretability), so the use of probabilistic scenarios is reserved for the main driver which is population. Clearly, given sufficient resources and appropriate models for the other socioeconomic drivers to build on, these factors can be treated in the same framework. This task is reserved for future research.
Concerning the task of quantifying food security risk, several indicators have been introduced in the literature so far. The UN have proposed the SDG Indicator 2.1.1 or otherwise the Prevalence of Undernourishment (PoU) indicator which is defined [^15] https://unstats.un.org/sdgs/metadata/?Text=&Goal=2&Target as ”an estimate of the proportion of the population whose habitual food consumption is insufficient to provide the dietary energy levels that are required to maintain a normal active and healthy life and is expressed as a percentage”. However, this indicator has been criticised on the grounds that its computation requires quite detailed data, e.g. periodic household surveys should be conducted, detailed information on food acquisition per household and habits, etc. Moreover, to correct potential biases, alternative data on the total amounts of available food for human consumption are needed. As a result, the computation of this indicator is quite a challenging task since rather detailed information is needed at household level, while forecasts concerning the index evolution either in the near future cannot be easily constructed.
Another famous relevant index is the Global Food Security Index (GFSI) [^16] https://impact.economist.com/sustainability/project/food-security-index/. The GFSI considers food affordability, availability, quality and safety, and sustainability and adaptation across 113 countries. The index is a dynamic quantitative and qualitative benchmark model constructed from 68 unique indicators that measure the drivers of food security across both developing and developed countries. Again, the estimation of such an index is far more complex than SDG 2.1.1, since its purpose is to express the state of food security at world (global) scale and unavoidably cannot be projected for prediction purposes.
The modelling approach discussed in the following sections aims at the construction of a reliable and robust with respect to model uncertainty) food security risk index, that can depict if a country is able to cover through its food system the minimum caloric requirements of its population. The index, should rely on data that are easy to obtain and from reliable sources, that are available for many countries over a sufficient depth in time. The proposed index is constructed in a way that can be projected to the future under the combined SSP-RCP scenarios taking into account each scenario’s effect to the underlying quantities that contribute to the indicator value. Note that dietary patterns, food habits and food affordability are not modelled in this version of the indicator.
@@ 3.1 Estimating minimum caloric requirements
The need for food is inelastic, in the sense that humans need a minimum and a maximum daily intake of calories for subsistence. The daily calories intake requirements vary per age group, sex, and lifestyle (e.g. level of activity) in a range of 1000 to 3200 calories depending on the category. In Tables 5 and 5 these requirements are shown, as proposed by the HHS/USDA [^17] United States Department of Health and Human Services and Department of Agriculture https://www.dietaryguidelines.gov/sites/default/files/2020-12/Dietary_Guidelines_for_Americans_2020-2025.pdf for the male and female population according to age group and activity level. Given the age structure of the male and female population (i.e. population pyramids per gender) for a country $c$, we may then obtain an estimate for the total daily recommended calories intake at year $t$, through the quantity
@@
where $a$ corresponds to the age groups mentioned in Tables 5 and 5, $P_{a,c,t}^{m},P_{a,c,t}^{f}$ denote the total male and female population for the respective age groups and $R_{a}^{m},R_{a}^{f}$ represent the calories requirements given in Tables 5 and 5. This estimate varies, depending on the activity level distribution of the population, however, one may obtain a lower bound for this quantity using the values for $R_{a}^{m},R_{a}^{f}$ for non active individuals or an upper bound using these values for the very active individuals. In this perspective, the quantity $C^{R}_{c,t}$ acts as a proxy for the total daily calorie needs (in terms of a lower or an upper estimate). Clearly, the current estimate may deviate from the actual daily minimum calorie requirements on account of malnutrition issues related to diet patterns, poverty or unequal income distribution, etc. However, as actual data for total calorie needs are not available, we consider $C^{R}_{c,t}$ as a reasonable proxy.
@@
Remark 3.1.
Minimum calories intake is modelled by the nominal value of the calories that are required by the population to survive, based only on the population structure and the nominal values for calories estimated by nutrition experts. Neither diet patterns or food product substitutions (e.g. when pasta is unavailable, replacing with rice, etc), or access to food are taken into account in this approach. The resulting estimate $C^{R}_{c,t}$ stated in (5) does not necessarily coincide either with the minimum nutritional needs or the food consumption of the country under study. Nutritional needs are subject to dietary patterns, available crops or food types. The estimation of minimum nutritional needs is a much more complex task than the estimation of the minimum calorie intake, since it relies on several aspects of food demand. Similarly, food consumption also involves aspects of food demand and supply (see e.g. Hasegawa et al. (2015); Valin et al. (2014); Van Dijk et al. (2021)). However, the estimate $C_{c,t}^{R}$ provided here, depends directly only on population structure of the country under study and does not take into account other factors, in an attempt to provide a rough, but as reliable as possible, estimate for the caloric requirements of the country’s population for subsistence. Clearly, a more detailed model could be considered as a next step in this modelling component, employing more detailed data concerning the nutrition patterns and habits, and the effects of active policies for increasing food affordability of the general population.

### 14 Dec 2023 v2

@@ 2.2 Shared Socioeconomic Pathways, Population and GDP Projections
The quantification of SSP scenarios using our methodology offers a notable advantage over traditional approaches, such as the UN methodology. In our approach, the scenario levels (i.e. sub-scenarios) are not preassigned; instead, they are endogenously determined by the historical data dynamics captured within the probabilistic population model. This model, inherently Bayesian, is calibrated using an extensive global database of past population data, which embeds significant historical insights into the population phenomenon [^4] The Bayesian model’s reliance on a vast global database infuses it with rich historical context, making it highly informative.. By applying this methodology to each trajectory in the sample for each country, we create three distinct sub-samples for both fertility rates and life expectancy. Each sub-sample corresponds to different potential realizations of the low, medium, and high intensity levels (sub-scenarios), providing critical statistical information like moments and variability within these scenarios. This approach offers a substantial advantage in the scenario-building process. It allows for the generation of more robust scenarios and the calculation of conditional expected values for other relevant quantities, based on the fundamental demographics modeled by these trajectories. A potential concern might be the representativeness of the Bayesian model’s generated sample of trajectories for key population drivers (TFR, $e_{0}$) and its capacity to encompass all plausible future states. However, given the model’s reliance on observed data from previous periods and its consideration of interconnections among countries and regions globally, it is reasonable to expect that any realistic case, reflective of the data available at the time of projection, can be simulated. Using a sufficiently large number of simulated trajectories, such as 100,000, should ensure reliability in the results. The specific country-level discrimination rules for each population driver are detailed in Table 2. It is important to note that migration levels could also be determined in a similar probabilistic manner (as suggested by Azose and Raftery (2015); Azose et al. (2016)). However, for the sake of simplicity and due to the minor role of migration in our application, we utilize pointwise net migration projections under each SSP scenario, as provided by the Wittgenstein Center database.
One can apply this methodology to a country of interest, thereby creating an extensive database of potential future population outcomes as derived from the population model (1). This database aligns with the various SSP scenarios and their respective sub-scenarios, focused on key population parameters. For instance, in a country categorized under the HiFert group (as per Table 1), the SSP2 scenario would encompass the TFR trajectories that fall within the Medium TFR sub-scenario, the e0 trajectories within the Medium e0 sub-scenario, and the net migration projections under the medium sub-scenario. These individual components – TFR, e0, and net migration – are integrated within the population model (1) to construct the comprehensive population trajectories that constitute the SSP2 scenario for this specific country. This probabilistic process, enables the calculation of various statistical measures, such as quantiles and moments for different population dimensions specific to each country. An example of the application of this methodology is illustrated in Figure 2, which displays the median population projections for Egypt (EGY) and Ethiopia (ETH) under each SSP scenario. Additionally, Figure 10 in the Appendix A presents the relative growth rates of these populations, along with the 90% uncertainty zones, for each scenario, while the instantaneous projected total population distributions at year 2050 for Egypt and Ethiopia under all SSP scenarios are illustrated in Appendix B.
@@
Building scenarios representing different pathways for the population and its age structure evolution, offers a vehicle for the estimation of key economic drivers, like labour force, gross domestic product (GDP) and others (see e.g. Aksoy et al. (2019)). In this paper such economic drivers scenarios (i.e. employing the various SSPs) are obtained using the global dynamic economic model MaGE Fouré et al. (2013); Fontagné et al. (2022). This model was developed by J. Fouré, A. Bénassy-Quéré and L. Fontagné and takes into account factors like aging populations, changes in fertility rates, and migration, and how these demographic shifts impact economic growth.  [^6] It is freely available from CEPII’s website [^5] http://www.cepii.fr/cepii/en/bdd_modele/presentation.asp?id=13 MaGE assumes that the world consists of economies of individual countries with each country $c$ characterized at time $t$ by a three-factor CES (Constant Elasticity of Substitution) production function with the capital and labour contributions modelled by the Cobb-Douglas parametric form
@@
where $I$ denotes the GDP, $K$ the capital, $L$ the labour force777child labour is not modelled by MaGE, $E$ the energy consumption with $c$ denoting the country and $t$ corresponding either to 1-year periods or 5-years periods. The elasticity parameters are assumed to be the same for all countries ($\alpha=0.31$, $\sigma=0.136$) and constant in time, while the parameters $A$ (Total Factor Productivity or TFP) and $B$ (Energy Productivity or Energy Efficiency888where energy efficiency is defined as an analogous quantity to the ratio between energy consumption and GDP) are assumed to be country specific and temporally varying. The model depends on population primarily through labor force and secondarily through the life cycle savings modelling which is introduced in the modelling of investment. MaGE allows the user to run generalized scenarios concerning the future states of the world economy, compatible with the SSP scenarios as far as population is concerned and produce relevant projections for each SSP scenario. In particular the model included in MaGE allows for the integration of demographics (including education) and economics to allow for predictions concerning population growth and economic quantities such as economic growth rates, GDP, energy consumption, etc.
@@ 2.3 A Combined Framework with Climate Scenarios
The impact of climate change on food security risk assessments has been a focal point in much of the relevant literature, as evidenced by studies like Schmidhuber and Tubiello (2007). Key discussions on the necessity of determining effective measures and actions to address food security challenges arising from climate change and extreme climate events are presented in Campbell et al. (2016) and Hasegawa et al. (2021). However, Hasegawa et al. (2018) suggest that the adverse effects of stringent climate mitigation policies could outweigh the direct impacts of climate change on global hunger and food consumption. Understanding how these climatic factors interact with population dynamics is crucial for anticipating future food security challenges. In the context of food security issues, this means that it important to consider critical environmental drivers such as temperature, precipitation, and water stress. These factors directly impact agricultural operations and, consequently, the overall capacity of the food system. To this end, the Representative Concentration Pathways (RCPs) Van Vuuren et al. (2011) provide essential scenarios. These scenarios project the increase in the Earth’s mean temperature by the year 2100, factoring in the implementation of various environmental policies. The RCPs are characterized by the degree of temperature increase and the extent of mitigation measures required in each scenario. For detailed specifications of each RCP scenario, please refer to Table 7 in the Appendix A.
@@
The current trend in developing realistic future world scenarios involves the integration of the Shared Socioeconomic Pathways (SSP) with the Representative Concentration Pathways (RCP). Initially, it might seem straightforward to conceive 30 distinct SSP-RCP scenarios, considering the ones outlined in Table 1 and the six standard RCP scenarios. However, the interaction between SSP and RCP scenarios is more complex, as they are interrelated despite focusing on different target quantities. Significant insights into these combined scenarios have been provided by the Coupled Model Intercomparison Project (CMIP) [^9] https://www.wcrp-climate.org/wgcm-cmip/wgcm-cmip6, which integrates diverse environmental models and utilizes extensive databases. The well-tested and available combined SSP-RCP scenarios, which include six distinct scenarios reflecting varying degrees of optimism about temperature change, are detailed in Table 3 Meinshausen et al. (2020).
In line with our objective to assess food security risks within a framework that aligns with both socioeconomic and climate pathways, we conduct our estimations based on these six combined scenarios. This involves developing appropriate models for food capacity, tailored to each scenario, as discussed in Section 3).
3 Towards a National Food Security Risk Index
The World Food Summit Summit (1996) defines food security as ”when all people, at all times, have physical and economic access to sufficient, safe and nutritious food that meets their dietary needs and food preferences for an active and healthy life”. According to this definition, food security encompasses four distinct dimensions [^10] https://www.worldbank.org/en/topic/agriculture/brief/food-security-update/what-is-food-security: (a) the physical availability of food, (b) the economic and physical access to food, (c) the utilization of food, and (d) the stability and sustainability of the other three dimensions over time.
In this context, we consider two primary metrics as the main components of food security risk: (a) the population’s minimum food requirements at a specific time point (focusing on caloric content rather than recommended dietary composition), and (b) the capacity of the national food system at the same time, quantified as the total caloric content of food available to the population as a result of the system’s functioning. It is important to note that our approach does not directly address finer aspects such as dietary needs, food habits, or poverty-related access issues at the household- or individual-level. [^11] Examples include Australia, Canada, the UK, and the USA, where high national food security contrasts with notable household-level food insecurity (see e.g. Loopstra (2018)). Nevertheless, the framework we propose is adaptable and can be extended to include these and other factors, should they be necessary for a more nuanced analysis. Henceforth, we use the term ’food requirements’ to refer specifically to caloric needs or intake.
Forecasting future caloric shortages is crucial for policymakers, enabling them to implement proactive strategies such as restructuring food production, optimizing land use, or planning imports. Developing effective policies requires a deep understanding of the dynamics influencing significant shifts in the food system. In this section, we introduce a streamlined statistical model that utilizes minimal and publicly available data to generate country-level, probabilistic projections of minimum caloric requirements and food system capacity, thereby quantifying food security risk. The model’s primary driver, population growth, is approached probabilistically, building upon the detailed model by Raftery et al. (2013) and its adaptations discussed in Section 2. While socioeconomic factors (like GDP and labor), natural resources (such as water resources and cropland area), and climate variables (temperature, precipitation) are integral to the model, they are not treated probabilistically. This limitation stems from constraints in computational resources, data availability, or the absence of suitable models, which would otherwise make the model excessively complex and challenging to interpret. Therefore, probabilistic scenarios are predominantly applied to population growth, the main driver. Expanding the probabilistic treatment to other drivers remains an objective for future research, contingent upon the availability of resources and appropriate models.
Our modelling approach is designed to create a food security risk index that is both reliable and robust against model uncertainties. This index assesses a country’s capacity to meet its population’s minimum caloric needs through the current function of its food system. The index’s construction is based on easily accessible, reliable data that span a significant historical period for numerous countries. It is tailored to project into the future under combined SSP-RCP scenarios, taking into account each scenario’s impact on the contributing factors to the index. It is important to note that in this version of the index, dietary patterns, food habits, and food affordability factors are not incorporated, with a focus instead on the core elements of food security risk assessment.
@@ 3.1 Estimating minimum caloric requirements
Minimum food requirements for subsistence, which are distinct from the caloric intake necessary for nutritional security [^12] It’s important to note that the bare minimum caloric intake for survival does not align with the calorie requirements ensuring nutritional security, are intrinsically linked to population size. These requirements are also influenced by the detailed age structure and activity levels of each age group. Humans have a fundamental requirement for a specific range of daily caloric intake, reflecting the inelastic nature of the need for food for subsistence This requirement varies across different age groups, genders, and lifestyles, notably influenced by levels of physical activity. Typically, the daily caloric intake necessary for an individual can range from approximately 1000 to 3200 calories, contingent upon these factors. In the provided Tables 5 and 5, we detail these caloric requirements. These tables reflect recommendations from the United States Department of Health and Human Services and Department of Agriculture (HHS/USDA) [^13] United States Department of Health and Human Services and Department of Agriculture https://www.dietaryguidelines.gov/sites/default/files/2020-12/Dietary_Guidelines_for_Americans_2020-2025.pdf, categorizing the necessary daily calorie intake for both male and female populations, segmented by age group and level of physical activity. Given the age structure of the male and female population (i.e. population pyramids per gender) for a country $c$, we may then obtain an estimate for the total daily recommended calories intake at year $t$, through the quantity
@@
where $a$ corresponds to the age groups mentioned in Tables 5 and 5, $P_{a,c,t}^{m},P_{a,c,t}^{f}$ denote the total male and female population for the respective age groups and $R_{a}^{m},R_{a}^{f}$ represent the calories requirements given in Tables 5 and 5. This estimate varies, depending on the activity level distribution of the population, however, one may obtain a lower bound for this quantity using the values for $R_{a}^{m},R_{a}^{f}$ for non-active individuals or an upper bound using these values for the very active individuals. In this perspective, the quantity $C^{R}_{c,t}$ acts as a proxy for the total daily calorie needs (in terms of a lower or an upper estimate). Clearly, the current estimate may deviate from the actual daily minimum calorie requirements on account of malnutrition issues related to diet patterns, poverty or unequal income distribution, etc. However, as actual data for total calorie needs are not available, we consider $C^{R}_{c,t}$ as a reasonable proxy.
Based on the above methodology, it becomes obvious that neither diet patterns or food product substitutions (e.g. when pasta is unavailable, replacing with rice, etc), or access to food are taken into account in this approach. The resulting estimate $C^{R}_{c,t}$ stated in (5) does not necessarily coincide either with the minimum nutritional needs or the food consumption of the country under study. Nutritional needs are subject to dietary patterns, available crops or food types. The estimation of minimum nutritional needs is a much more complex task than the estimation of the minimum calorie intake, since it relies on several aspects of food demand. Similarly, food consumption also involves aspects of food demand and supply (see e.g. Hasegawa et al. (2015); Valin et al. (2014); Van Dijk et al. (2021)). However, the estimate $C_{c,t}^{R}$ provided here, depends directly only on population structure of the country under study and does not take into account other factors, in an attempt to provide a rough, but as reliable as possible, estimate for the caloric requirements of the country’s population for subsistence. Clearly, a more detailed model could be considered as a next step in this modelling component, employing more detailed data concerning the nutrition patterns and habits, and the effects of active policies for increasing food affordability of the general population.
Utilizing the outlined methodology, it is important to acknowledge that this approach does not factor in dietary patterns, food product substitutions (such as replacing pasta with rice when the former is unavailable), or accessibility to food. The estimated total caloric requirement $C^{R}_{c,t}$, as detailed in (5), is not synonymous with either the minimum nutritional needs or the actual food consumption of the country being studied. Nutritional requirements are influenced by factors like dietary habits, availability of various crops, and types of food. Estimating minimum nutritional needs is inherently more complex than calculating minimum calorie intake, as it entails a broader array of food consumption aspects, Similarly, such as demand and supply elements in the food system, as discussed in Hasegawa et al. (2015); Valin et al. (2014); Van Dijk et al. (2021). However, the estimate $C_{c,t}^{R}$ we provide is based solely on the population structure of the country in question, deliberately excluding other variables. This approach aims to offer a basic yet reliable estimation of the population’s caloric requirements for subsistence. It’s clear that a more comprehensive model, incorporating detailed data on nutrition patterns, eating habits, and the impact of policies aimed at improving food affordability for the general population, could be developed as a subsequent enhancement to this component of the modeling process.